Use a triple integral to compute the volume of the region bounded by curves y = 2-2x, x = 0,, and y=0 in the xy plane and the surface defined above by z = x^2
Answer #1

OLZ 2x2 04x427) oty 22 2 2 / x² dz dedy 0 x 0 ko x=24 (x²) dady 4- o a=0r = 3 / (ej to the Playsidy